Determine the length of the curve 𝑥 = 𝑦^2 /2 for 0 ≤ 𝑥 ≤ 1/2 . Assume 𝑦 positive.




1
Expert's answer
2021-12-13T17:12:41-0500

length of the curve:


"L=\int\sqrt{1+(y')^2}dx"


"y=\sqrt{2x}"

"y'=1/(\sqrt{2x})"


"L=\int^{1/2}_0\sqrt{1+1/(2x)}dx=(x\sqrt{1+1/(2x)}+\frac{ln(\sqrt{2+1/x}+\sqrt2)-ln(\sqrt{2+1/x}-\sqrt2))}{4}|^{1/2}_0="


"=1/\sqrt 2+\frac{ln(\sqrt2+1)-ln(\sqrt2-1)}{4}"